# In the Eye of MLLM: Benchmarking Egocentric Video Intent Understanding with Gaze-Guided Prompting

### v1

@@ Abstract
The emergence of advanced multimodal large language models (MLLMs) has significantly enhanced AI assistants’ ability to process complex information across modalities. Recently, egocentric videos, by directly capturing user focus, actions, and context in an unified coordinate, offer an exciting opportunity to enable proactive and personalized AI user experiences with MLLMs. However, existing benchmarks overlook the crucial role of gaze as an indicator of user intent. To address this gap, we introduce EgoGazeVQA, an egocentric gaze-guided video question answering benchmark that leverages gaze information to improve the understanding of longer daily-life videos. EgoGazeVQA consists of gaze-based QA pairs generated by MLLMs and refined by human annotators. Our experiments reveal that existing MLLMs struggle to accurately interpret user intentions. In contrast, our gaze-guided intent prompting methods significantly enhance performance by integrating spatial, temporal, and intent-related cues. We further conduct experiments on gaze-related fine-tuning and analyze how gaze estimation accuracy impacts prompting effectiveness. These results underscore the value of gaze for more personalized and effective AI assistants in egocentric settings.
@@ 1 Introduction
The rapid advancement of multimodal large language models (MLLMs)hurst2024gpt ; team2024gemini ; chen2024internvl ; wang2024qwen2 has enabled highly intelligent AI assistants capable of understanding and generating complex information across modalities. However, users often have to design intricate prompts to effectively convey their intentions and personalize the assistant. This challenge raises an essential question: Can existing MLLM enable a more proactive AI user experience– one where the assistant seamlessly analyzes the users’ underlying intention and delivers personalized content without heavily replying on complex prompting?
Egocentric videos may serve as an ideal solution to this challenge: the unique viewpoint aligns human visual perception, attention, action, and surrounding scene-context within the same egocentric coordinate system, and thereby offers valuable insights into an individual’s experiences, actions, and intentions. Despite the potential of egocentric capture to pave the way for a more personalized and proactive AI user experience konrad2024gazegpt , existing MLLMs have been predominantly developed and evaluated using exocentric video benchmarks yu2019activitynet ; zhou2025egotextvqa ; li2023intentqa ; fu2024video ; wang2024lvbench ; li2024mvbench ; cores2024tvbench ; wu2024star . Recent works introduced several egocentric VQA benchmarks for Episodic Memory barmann2022did , Long-Form Video reasoning shen2018egocentric , Scene Layout analysis huang2025building , and Scene-Text understanding zhou2025egotextvqa . However, these benchmarks have overlooked a crucial egocentric cue—Gaze—which provides direct insights into user attention and intent. For egocentric AI assistant, the majority of user questions are likely to be inherently conditioned on what the user is looking at. Taking the examples in Figure 1, the MLLMs may run into failure mode when addressing questions about user fixation without gaze signals. To bridge this gap, we introduce a timely benchmark – EgoGazeVQA, an Egocentric Gaze-Guided Video Question Answering Benchmark, specifically designed to assess whether MLLMs can utilize egocentric gaze signals for analyzing daily-routine videos.
@@
    We present empirical analysis demonstrating how LoRA-based fine-tuning can enhances a model’s ability to leverage gaze signals for understanding human intentions from an egocentric perspective.
@@ 4.2 Gaze-guided intent prompting
To investigate whether incorporating 2D gaze coordinates enhances a model’s understanding of human intention on EgoGazeVQA benchmark, we explore three prompting strategies, as illustrated in Figure 3. Each strategy varies in how gaze information is encoded and presented to the MLLMs, allowing us to examine different levels of visual and textual alignment. The detailed inference prompt template can be found in the Suppl.
@@
Sequential gaze salience map as prompt (GazeS). Considering the strong temporal correlation between the video context and gaze trajectory, we further model gaze over time using salience maps. Specifically, we progressively enhance the intensities of gaze saliency maps as the video advances and consolidate them into a single saliency map that captures both spatial and temporal intentions. Regions revisited by the user’s gaze are assigned higher intensities, effectively mimicking human visual scanning patterns over time. Pseudocode for generating these salience maps is provided in Algorithm 1 in Appendix B.
@@ 4.3 Results on EgoGazeVQA
Qwen2.5-VL Qwen2.5-7B 48.8 $\uparrow$1.7 50.7 $\uparrow$4.8 63.9 $\uparrow$8.3 56.5 $\uparrow$2.8 46.7 $\uparrow$3.8 49.7 $\uparrow$3.7
Qwen2.5-VL Qwen2.5-72B 58.5 $\uparrow$2.5 60.9 $\uparrow$8.1 70.8 $\uparrow$1.4 69.4 $\uparrow$2.8 59.9 $\uparrow$4.9 60.5 $\uparrow$4.9
@@
Interestingly, our benchmark also presents challenges to human subjects; however, humans consistently outperform the best-performing model (Qwen2.5-VL-72B) across all categories, particularly in spatial and temporal reasoning. This indicates that the task is limited by current model capabilities. The sizable gap indicates the opportunity for future work to better leverage gaze cues on intention understanding.
@@
Qwen2.5-VL Qwen2.5-7B 49.6 $\uparrow$1.3 55.7 $\uparrow$4.2 66.1 $\uparrow$3.4 55.6 $\uparrow$3.7 50.7 $\uparrow$1.5 36.7 $\uparrow$8.5
Qwen2.5-VL Qwen2.5-72B 58.7 $\uparrow$2.1 65.3 $\uparrow$10.5 71.2 $\uparrow$3.4 69.4 $\uparrow$7.5 52.3 $\uparrow$2.8 52.1 $\uparrow$6.4
@@ 4.5 Additional analsyis
Table 7: Prompt-based gaze estiamtion with MLLM and effect on EgoGazeVQA. GT: Ground Truth. GS: Gaze Estimation
  Datasets   Qwen2.5 VL-72B  Acc   MSE
  w/o   GT   $\Delta(\text{GT}-\text{w/o})$   GE   $\Delta(\text{GE}-\text{w/o})$
@@
Error Type Spatial Temporal Causal
Inaccurate 17.89 21.03 8.05
Irrelevant 16.78 13.87 10.96
Unanswerable 15.88 9.84 6.26
Table 9: Effect of varying numbers of input frames
Frames Spatial Temporal Causal Average
4 46.9 45.5 78.6 57.0
6 47.4 45.0 79.0 57.1
[…]
14 49.2 43.3 79.0 57.1
18 47.2 42.6 79.2 56.3
Prompted MLLM gaze estimation and impact on EgoGazeVQA. To further show that explicit gaze signal is vital for MLLM to understand human intent, we further query Qwen2.5-VL-72B for egocentric gaze estimation using the following prompt“Mark the coordinates the camera wearer is looking at""; the returned $(x,y)$ forms a frame-wise gaze map that is fed back as a sequential gaze salience map prompt (GazeS) for EgoGazeVQA. Table 7 shows that gaze estimation accuracy is strongly correlated with performance improvement. (for example, MSE = 0.038 in EgoExo), the evaluation results lags slightly behind those obtained using the ground-truth gaze. In contrast, when gaze estimation is less accurate, such as in EGTEA (MSE = 0.119), the benefit from GE diminishes drastically and may even slightly hurt performance. These results suggest that our proposed method is capable of effectively leveraging gaze signals—when they are reliable—to improve question answering.
Effect of varying numbers of input frames. Since MLLMs perceive video through sparsely sampled images, we vary the number of uniformly sampled frames fed to Qwen2.5-VL-72B and report the results in Table 9. Notably, adopting 9 frames as the default setting offers the best trade-off between accuracy and computational cost.
@@ 5 Conclusion
In this paper, we introduced EgoGazeVQA, an egocentric gaze-guided video question answering benchmark designed to evaluate the ability of existing MLLMs to interpret user intentions from first-person videos. Additionally, we proposed and examined gaze-guided prompting methods that utilize gaze information as a direct indicator of user focus and intent, enabling a deeper understanding of spatial, temporal, and causal relationships in egocentric contexts. Our experiments show that integrating gaze cues significantly enhances MLLM performance across various scenarios and activities. We believe that EgoGazeVQA represents an important step towards developing proactive and personalized AI assistants with egocentric MLLMs. Furthermore, our work opens up promising directions for research on multi-sensory MLLMs and MLLMs explainability.
@@ Appendix A Appendix Overview
Code Repository https://github.com/anonyupload/EgoGazeVQA Apache-2.0 license
Data https://huggingface.co/datasets/anonupload/EgoGazeVQA-91-nips25DB CC BY 4.0
@@ Appendix C More experiments results
C.1 Baseline Comparison with CLIP, EgoVLP, and BLIP-2
In Table 3 of the main paper, we report performance of representative MLLMs on EgoGazeVQA to establish demonstrative baselines and highlight the significant room for improvement on gaze-relevant tasks. Here, we provide results using CLIP (ViT-H-14) and EgoVLP as simple baselines, where we compute the cosine similarity between the visual embeddings of the input and the text embeddings of each answer option, selecting the one with the highest similarity as the prediction. We also evaluate BLIP-2, which generates answers based on the visual embeddings and input question. We calculate the cross-entropy loss between the generated answer and each of the candidate options, and select the one with the lowest loss as the most likely match.
For both baselines, we follow the same video preprocessing pipeline described in the main paper. To mimic human visual attention, we use gaze fixation as an attentional map to select around 20% of each frame, resulting in what we term the GazeMap variant (i.e., CLIP-GazeMap, EgoVLP-GazeMap, and BLIP-2-GazeMap), with all other components unchanged.
Table 10: Performance comparison of simple baseline models and Qwen2.5-VL-72B variants on EgoGazeVQA.
  Method Similarity Performance
Human N/A 76.8
CLIP ViT-H-14 Video-A 21.63
CLIP-GazeMap Video-A 21.79
EgoVLP Video-A 22.95
EgoVLP-GazeMap Video-A 22.65
BLIP-2 N/A 27.60
BLIP-2-GazeMap N/A 27.21
Qwen2.5-VL-72B N/A 60.5
Qwen2.5-VL-72B-GazeVisual N/A 63.9
Notably, these simple baselines perform only marginally better than random guessing and fall far short of human expert performance, indicating that CLIP-based methods struggle to tackle this challenging benchmark. Moreover, simply narrowing the visual field through gaze cropping proves insufficient for fully leveraging gaze information with existing vision encoders. This reinforces the need for the specialized prompting strategies proposed in our work, which incorporate spatiotemporal and intent-aware cues.
C.2 The complexity of the task purely based on number of targets in the scene
We run YOLOv11 on the entire test set to estimate how many discernible targets each clip contains. The average statistics are summarized below:
Table 11: Scene complexity and Qwen2.5VL-72B performance by location.
  By Location Kitchen Living Room Medical Area Garage Others
objects classes 4.1 5.1 3.9 2.9 4.9
num objects 7.1 8.3 7.7 5.2 7.7
Qwen2.5VL-72B (acc) 58.5 60.9 70.8 69.4 59.9
Taking Qwen2.5VL-72B as an example, we observe that although the Living Room contains more objects than the Kitchen, model performance is similar. Likewise, the Medical Area has a comparable number of objects to the Kitchen, yet achieves significantly higher accuracy.
Table 12: Scene complexity and Qwen2.5VL-72B performance by activity.
  By Activity Cooking Gaming Health Mechanics Creative Others
objects classes 4.1 5.4 3.8 2.8 3.5 5.1
num objects 7.1 8.7 7.4 5.2 4.8 8.4
Qwen2.5VL-72B (acc) 58.7 65.3 71.2 69.4 52.3 52.1
In activity-based scenes, the Gaming contains the most objects on average, but the model performs better than in Cooking, which has fewer targets.
These findings suggest that benchmark complexity is not solely determined by the number of visual targets. Instead, it is also shaped by the location and activity context, where different spatial structure, temporal dependencies, and the implicit goals of the human subject all affect how the model interprets human intention.
C.3 Analysis of out-of-distribution (OOD) performance.
Most of the results in our main experiments are based on gaze-guided prompting without any parameter tuning. Since the model weights are frozen throughout our model, the generalization on other tasks or benchmarks won’t be affected by our method.
Gaze-Guided Prompting approach (main results): Our main method is the intent-guided prompting, where the model weights are frozen and gaze is incorporated purely via prompting. In this training-free setting, the original model’s general capabilities remain unchanged, while it gains the ability to answer gaze-related queries.
Lora Fine-tuning experiments (analytic purpose): The LoRA fine-tuning is intended as an analytic experiment — to study potential overfitting and to encourage researchers to explore more effective training methods for addressing EgoGazeVQA without overfitting — rather than as the optimal solution to our task. In fact, how to design a training pipeline that boosts performance on EgoGazeVQA without losing generalization ability on other tasks is an important open challenge we aim to introduce to the community for future exploration. We will clarify this in the revision.
Additional Generalization experiments: Though solving the generalization problem is beyond the scope of this benchmark paper, we provide an initial attempt to optimize the LoRA SFT training data to better balance performance on EgoGazeVQA and EgoSchema. Instead of training solely on video–question–answer pairs, which risks overfitting, we concatenated the text-based gaze-guided prompt (GazeT) with each question, enabling the model to learn meaningful knowledge for gaze-specific associations while reducing overfitting on general questions.
We applied this to Qwen2.5-VL-7B, and tested the models on EgoSchema for OOD generalization. Note that all LoRA SFT models are trained on EGTEA tested on Ego4D and EgoExo. The results are shown below:
Table 13: Comparison of SFT training strategies on EgoSchema and EgoGazeVQA benchmarks.
  Methods SFT Data Question Format EgoSchema Avg (%) EgoGazeVQA Avg (%)
Zero-shot 7B N/A 60.72 50.26
LoRA SFT 7B Question only 48.30 (-12.42) 68.85 (+18.59)
LoRA SFT 7B Textual Gaze Prompt + Question 58.32 (-2.40) 67.00 (+16.74)
For fair comparison, all above experiments from the above table are using textual gaze prompt (GazeV) for EgoGazeVQA. From the 7B model results, SFT using gaze-guided prompts together with questions helps mitigate overfitting compared with question-only training, achieving performance close to the original model on EgoSchema while also delivering notable improvements on EgoGazeVQA. Here, we focus on examining this data optimization strategy with textual gaze prompts (GazeT), and we will explore applying similar strategies to visual gaze prompt (GazeV) and saliency gaze prompt (GazeS) in future work.
C.4 Full fine-tuning results compared with LoRA fine-tuning
We conducted full fine-tuning on Qwen2.5VL-7B using both the EGTEA split and the Ego4D + EgoExo datasets. The results are presented below:
Table 14: Comparison between full fine-tuning and LoRA fine-tuning on Qwen2.5VL-7B.
  Methods Training Data Test Data Spatial Temporal Causal Avg
Zero-shot N/A Ego4D + EgoExo 40.4 43.3 78.2 54.0
LoRA SFT EGTEA Ego4D + EgoExo 67.7 56.5 84.4 69.5
Full SFT EGTEA Ego4D + EgoExo 69.6 58.9 82.7 70.4
Zero-shot N/A EGTEA 45.7 32.9 77.3 52.0
LoRA SFT Ego4D + EgoExo EGTEA 66.5 47.0 82.8 65.4
Full SFT Ego4D + EgoExo EGTEA 69.5 48.2 77.9 65.2
As shown, the performance improvement from full fine-tuning is comparable to LoRA, while requiring significantly more trainable parameters. This indicates parameter-efficient LoRA is still a competitive and practical choice.
C.5 Analysis of gaze estimation errors impact model performance in specific scenarios
We break down the results by reasoning type: spatial, temporal, and causal. As shown in the table below (GT refers to using gaze ground truth in our prompting strategy, while GE refers to using gaze estimation.):
Table 15: Impact of gaze estimation (GE) versus ground truth gaze (GT) on spatial, temporal, and causal reasoning performance.
  Category w/o (baseline) GT $\Delta$(GT - w/o) GE $\Delta$(GE - w/o)
Spatial 57.1 64.3 +7.2 53.8 -3.3
Temporal 45.2 50.3 +5.1 48.6 +3.4
Causal 79.4 84.3 +4.9 82.2 +2.8
From the results and visualizations, we can draw the following conclusions:
We observe that gaze-guided prompting using gaze estimation leads to performance degradation in spatial questions. This suggests that spatial reasoning is highly sensitive to gaze estimation accuracy, as it often depends on precise spatial grounding. For instance, if the estimated gaze falls slightly off the object being manipulated—especially when nearby objects are close together, such as mistaking a knife for a cutting board—the model may misinterpret the target of the action and produce an incorrect response.
In contrast, temporal questions rely more on in-context temporal cues and are less affected by gaze estimation noise. Still, if the estimated gaze lands on a static background object rather than a newly appearing item, the model may miss key temporal transitions.
For causal reasoning, if the gaze is misestimated to a nearby object rather than the causal trigger (e.g., glancing at a mug instead of the hand turning on a kettle), the model might fail to infer the correct cause-effect relationship.
Why we report MSE for prompt-based gaze estiamtion. Most egocentric gaze-estimationlai2022eye huang2018predicting output a heat-map and therefore adopt hit/miss metrics such as F1 or AUC within a fixed angular tolerance. In contrast, our prompt returns an explicit $(x,y)$ coordinate, not a density map. For a single point prediction the most natural error measure is the Euclidean distance to the ground-truth fixation, which—after normalising image size—reduces to a mean-squared error (MSE). Using MSE avoids choosing an arbitrary threshold, is scale-independent, and directly reflects how far the predicted fixation deviates from the true one; hence we employ it in Table 7.

### v2

@@ Abstract
The emergence of advanced multimodal large language models (MLLMs) has significantly enhanced AI assistants’ ability to process complex information across modalities. Recently, egocentric videos, by directly capturing user focus, actions, and context in an unified coordinate, offer an exciting opportunity to enable proactive and personalized AI user experiences with MLLMs. However, existing benchmarks overlook the crucial role of gaze as an indicator of user intent. To address this gap, we introduce EgoGazeVQA, an egocentric gaze-guided video question answering benchmark that leverages gaze information to improve the understanding of longer daily-life videos. EgoGazeVQA consists of gaze-based QA pairs generated by MLLMs and refined by human annotators. Our experiments reveal that existing MLLMs struggle to accurately interpret user intentions. In contrast, our gaze-guided intent prompting methods significantly enhance performance by integrating spatial, temporal, and intent-related cues. We further conduct experiments on gaze-related fine-tuning and analyze how gaze estimation accuracy impacts prompting effectiveness. These results underscore the value of gaze for more personalized and effective AI assistants in egocentric settings. Project page: https://taiyi98.github.io/projects/EgoGazeVQA.
@@ 1 Introduction
The rapid advancement of multimodal large language models (MLLMs) hurst2024gpt ; team2024gemini ; chen2024internvl ; wang2024qwen2 has enabled highly intelligent AI assistants capable of understanding and generating complex information across modalities. This capability is not only valuable for enhancing current AI systems but also crucial for advancing future human–AI collaboration systems, thereby enabling new AI paradigms such as cobodied AI  lu2025cobodied . However, to effectively collaborate with users, existing systems often require them to design intricate prompts to convey their intentions and personalize the assistant’s responses. This challenge raises an essential question: Can existing MLLM enable a more proactive AI user experience– one where the assistant seamlessly analyzes the users’ underlying intention and delivers personalized content without heavily replying on complex prompting?
Egocentric videos may serve as an ideal solution to this challenge: the unique viewpoint aligns human visual perception, attention, action, and surrounding scene-context within the same egocentric coordinate system, and thereby offers valuable insights into an individual’s experiences, actions, and intentions. Despite the potential of egocentric capture to pave the way for a more personalized and proactive AI user experience konrad2024gazegpt , existing MLLMs have been predominantly developed and evaluated using exocentric video benchmarks yu2019activitynet ; zhou2025egotextvqa ; li2023intentqa ; fu2024video ; wang2024lvbench ; li2024mvbench ; cores2024tvbench ; wu2024star . Recent works introduced several egocentric VQA benchmarks for Episodic Memory barmann2022did , Long-Form Video reasoning shen2018egocentric , Scene Layout analysis huang2025building , and Scene-Text understanding zhou2025egotextvqa . However, these benchmarks have overlooked a crucial egocentric cue—Gaze—which provides direct insights into user attention bao2025gazegene and intent. For egocentric AI assistant, the majority of user questions are likely to be inherently conditioned on what the user is looking at. Taking the examples in Figure 1, the MLLMs may run into failure mode when addressing questions about user fixation without gaze signals. To bridge this gap, we introduce a timely benchmark – EgoGazeVQA, an Egocentric Gaze-Guided Video Question Answering Benchmark, specifically designed to assess whether MLLMs can utilize egocentric gaze signals for analyzing daily-routine videos.
@@
    We present empirical analysis demonstrating how fine-tuning can enhances a model’s ability to leverage gaze signals for understanding human intentions from an egocentric perspective.
@@ 4.2 Gaze-guided intent prompting
To investigate whether incorporating 2D gaze coordinates enhances a model’s understanding of human intention on EgoGazeVQA benchmark, we explore three prompting strategies, as illustrated in Figure 3. Each strategy varies in how gaze information is encoded and presented to the MLLMs, allowing us to examine different levels of visual and textual alignment. The detailed inference prompt template can be found in Appendix E.
@@
Sequential gaze salience map as prompt (GazeS). Considering the strong temporal correlation between the video context and gaze trajectory, we further model gaze over time using salience maps. Specifically, we progressively enhance the intensities of gaze saliency maps as the video advances and consolidate them into a single saliency map that captures both spatial and temporal intentions. Regions revisited by the user’s gaze are assigned higher intensities, effectively mimicking human visual scanning patterns over time. Pseudocode for generating these salience maps is provided in Algorithm 1 of Appendix B.
4.3 Baseline Comparison with CLIP, EgoVLP, and BLIP-2
In Table 3, we report performance of representative MLLMs on EgoGazeVQA to establish demonstrative baselines and highlight the significant room for improvement on gaze-relevant tasks. In Table 4, we provide results using CLIP (ViT-H-14) and EgoVLP as simple baselines, where we compute the cosine similarity between the visual embeddings of the input and the text embeddings of each answer option, selecting the one with the highest similarity as the prediction. We also evaluate BLIP-2, which generates answers based on the visual embeddings and input question. We calculate the cross-entropy loss between the generated answer and each of the candidate options, and select the one with the lowest loss as the most likely match.
For both baselines, we follow the same video preprocessing pipeline described in the main paper. To mimic human visual attention, we use gaze fixation as an attentional map to select around 20% of each frame, resulting in what we term the GazeMap variant .
Table 4: Performance comparison of simple baseline models and Qwen2.5-VL-72B variants (i.e., CLIP-GazeMap, EgoVLP-GazeMap, and BLIP-2-GazeMap) on EgoGazeVQA.
     Method    Similarity    Performance
   CLIP ViT-H-14    Video-A    21.63
   CLIP-GazeMap    Video-A    21.79
   EgoVLP    Video-A    22.95
   EgoVLP-GazeMap    Video-A    22.65
   BLIP-2    N/A    27.60
   BLIP-2-GazeMap    N/A    27.21
   Qwen2.5-VL-72B    N/A    60.5
   Qwen2.5-VL-72B-GazeVisual    N/A    63.9
Notably, these simple baselines perform only marginally better than random guessing and fall far short of human expert performance, indicating that CLIP-based methods struggle to tackle this challenging benchmark. Moreover, simply narrowing the visual field through gaze cropping proves insufficient for fully leveraging gaze information with existing vision encoders.
@@ 4.4 Results on EgoGazeVQA
Qwen2.5-VL-7B Qwen2.5-7B 48.8 $\uparrow$1.7 50.7 $\uparrow$4.8 63.9 $\uparrow$8.3 56.5 $\uparrow$2.8 46.7 $\uparrow$3.8 49.7 $\uparrow$3.7
Qwen2.5-VL-72B Qwen2.5-72B 58.5 $\uparrow$2.5 60.9 $\uparrow$8.1 70.8 $\uparrow$1.4 69.4 $\uparrow$2.8 59.9 $\uparrow$4.9 60.5 $\uparrow$4.9
@@
Our benchmark also presents challenges to human subjects. However, humans consistently outperform the best-performing model (Qwen2.5-VL-72B) across all categories, particularly in spatial and temporal reasoning. This indicates that the task is limited by current model capabilities. The sizable gap indicates the opportunity for future work to better leverage gaze cues on intention understanding. We also vary the number of uniformly sampled frames fed to Qwen2.5-VL-72B and report the results in Appendix C.
     Error Type    Spatial    Temporal    Causal
   Inaccurate    17.89    21.03    8.05
   Irrelevant    16.78    13.87    10.96
   Unanswerable    15.88    9.84    6.26
@@
Qwen2.5-VL-7B Qwen2.5-7B 49.6 $\uparrow$1.3 55.7 $\uparrow$4.2 66.1 $\uparrow$3.4 55.6 $\uparrow$3.7 50.7 $\uparrow$1.5 36.7 $\uparrow$8.5
Qwen2.5-VL-72B Qwen2.5-72B 58.7 $\uparrow$2.1 65.3 $\uparrow$10.5 71.2 $\uparrow$3.4 69.4 $\uparrow$7.5 52.3 $\uparrow$2.8 52.1 $\uparrow$6.4
@@ 4.6 Additional analsyis
Table 9: Prompt-based gaze estiamtion with MLLM (Qwen2.5 VL-72B) and effect on EgoGazeVQA. GT: Ground Truth. GS: Gaze Estimation.
    Datasets   w/o   GT   $\Delta(\text{GT}-\text{w/o})$   GE   $\Delta(\text{GE}-\text{w/o})$   MSE
@@
Prompted MLLM gaze estimation and impact on EgoGazeVQA. To further show that explicit gaze signal is vital for MLLM to understand human intent, we further query Qwen2.5-VL-72B for egocentric gaze estimation using the following prompt“Mark the coordinates the camera wearer is looking at""; the returned $(x,y)$ forms a frame-wise gaze map that is fed back as a sequential gaze salience map prompt (GazeS) for EgoGazeVQA.
Table 9 shows that gaze estimation accuracy is strongly correlated with performance improvement. (for example, MSE = 0.038 in EgoExo), the evaluation results lags slightly behind those obtained using the ground-truth gaze. In contrast, when gaze estimation is less accurate, the benefit from GE diminishes drastically and may even slightly hurt performance.
@@ 5 Conclusion
In this paper, we introduced EgoGazeVQA, an egocentric gaze-guided video question answering benchmark designed to evaluate the ability of existing MLLMs to interpret user intentions from first-person videos. Additionally, we proposed and examined gaze-guided prompting methods that utilize gaze information as a direct indicator of user focus and intent, enabling a deeper understanding of spatial, temporal, and causal relationships in egocentric contexts. Our experiments show that integrating gaze cues significantly enhances MLLM performance across various scenarios and activities. We believe that EgoGazeVQA represents an important step towards developing proactive and personalized AI assistants with egocentric MLLMs. Furthermore, our work opens up promising directions for research on multi-sensory MLLMs and MLLMs explainability. Moreover, by grounding MLLMs in human-perspective signals like gaze, our work also provide foundations for cobodied intelligence.
@@ Appendix A Appendix Overview
Code Repository https://github.com/taiyi98/EgoGazeVQA Apache-2.0 license
Data https://huggingface.co/datasets/taiyi09/EgoGazeVQA CC BY 4.0
@@ Appendix C More experiments results
C.1 Effect of varying numbers of input frames.
Since MLLMs perceive video through sparsely sampled images, we vary the number of uniformly sampled frames fed to Qwen2.5-VL-72B and report the results in Table 10. Notably, adopting 9 frames as the default setting offers the best trade-off between accuracy and computational cost.
Table 10: Effect of varying numbers of input frames
    Frames   Spatial   Temporal   Causal   Average
  4   46.9   45.5   78.6   57.0
  6   47.4   45.0   79.0   57.1
[…]
  14   49.2   43.3   79.0   57.1
  18   47.2   42.6   79.2   56.3
C.2 Why we report MSE for prompt-based gaze estiamtion.
Most egocentric gaze-estimationlai2022eye huang2018predicting output a heat-map and therefore adopt hit/miss metrics such as F1 or AUC within a fixed angular tolerance. In contrast, our prompt returns an explicit $(x,y)$ coordinate, not a density map. For a single point prediction the most natural error measure is the Euclidean distance to the ground-truth fixation, which—after normalising image size—reduces to a mean-squared error (MSE). Using MSE avoids choosing an arbitrary threshold, is scale-independent, and directly reflects how far the predicted fixation deviates from the true one; hence we employ it in Table 9.